OpenStudy (anonymous):

solve a squared-2a-3/a squared-9a +18 minus a squared -5a-6/a squared+9a+8

OpenStudy (magbak):

do not listen to cambel he gives wrong answers.

OpenStudy (campbell_st):

well your question looks like \[\frac{a^2 -2a -3}{a^2 - 9a + 18} - \frac{a^2 -5a -6}{a^2 + 9a + 8}\] your 1st step should be to factor all the quadratics. \[\frac{(a +1)(a -3)}{(a - 3)(a-6)} - \frac{(a - 6)(a + 1)}{(a + 1)(a + 8)}\] cancel the common factors... then get the common denominator, and cross multiply. then distribute and collect like terms and simplify.

OpenStudy (anonymous):

Thanks! Greatly appreciated. I came up with 2asquared-28/(a-6)(a+8). Hope its right. Thanks

OpenStudy (campbell_st):

ummm I think the denominator is correct but the numerator may be \[\frac{a^2 + 9a + 8 - a^2 + 12a -36}{(x -6)(a +8)} = \frac{21a - 28}{(a - 6)(a + 8)}\] and the numerator can be factored to 7(3x - 4)

OpenStudy (anonymous):

Thanks for breaking it down. HUGE help. Thanks again